The Saunders Company is trying to do some planning for the coming month. The following information was gathered from the last month.

Sales revenue $300,000
Fixed manufacturing expenses    110,000
Fixed marketing expenses 20,000
Fixed administrative expenses 18,500
Sales price $20/unit
Variable manufacturing costs $7/unit
Variable marketing costs $2/unit

If Saunders wants to earn operating income of $44,000 next month, how many units must be sold?
Answer #1

Calculate required unit Sales :

Fixed Cost = 110000+20000+18500 = $148500

Variable cost per unit = 7+2 = $9 per unit

Required unit sales = (Fixed cost+desired profit)/Sale price per unit-Variable cost per unit

= (148500+44000)/(20-9)

Required unit sales = 17500 Units